OpenStudy (anonymous):

Julio wants to deposit $1500 into savings accounts at three different banks: Bank of Chance, Merchant Bank, and Utopian Financing. He will deposit five times as much into Merchant Bank as Bank of Chance because they offer a higher interest rate. He also expects the Utopian Financing deposit to be only 25% of the total of the other two deposits. How much will Julio deposit into the Utopian Financing savings account?

OpenStudy (anonymous):

C + 5c + 6C/4 = $1,500 --> c = $1,500 / 7.5 = $200 on Bank of Chance => 5 * 200 = $ 1,000 on Merchant Bank => Thus Utopian = 1,200/ 4 = $300